pls. help me solve this problem
thank you.
prob 1. A boat goes 6 miles downstream in a river in 30 minutes and returns upstream in 45 minutes. Find the speed of the boat in still water and the speed of the current.

Answer by mananth(16946)

still water speed = xmph
current speed = y mph
...
Time = distance / speed
upstream time = 6/ (x-y ) =(3/4)
3(x-y ) =6*4
x-y = 8.................1
downstream speed = 6/(x+y) = (1/2)
x+y = 12.................2
add 1 & 2
2x=20
x=10 mph speed of boat.
x-y =8
so y =2 mph speed of current